OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)? y = 2x – 8 y = – 2x + 9 y = 2x – 11 y = – 2x + 6

OpenStudy (anonymous):

so lets find the gradient using the two points given. Gradient/slope formula is \[m=\frac{y2-y1}{x2-x1}\] so \[m=\frac{-1--7}{5-2}=\frac{-1+7}{3}=\frac{6}{3}=2\] using the gradient and one point given, lets use (5,-1) and the formula to find a straight line which is y2-y1 = m (x2-x1) y - (-1) = 2 (x-5) y + 1 = 2x-10 y = 2x - 10 -1 y = 2x - 11 so there we have C as the answer.

OpenStudy (anonymous):

Alright Thank you!!!!!!! :)